OpenStudy (xoterrixo13):

find the midpoint of the segment wth the given end points d(1,2) and e(-3,6)

OpenStudy (tyteen4a03):

The midpoint formula is: x = (x1+x2)/2, y = (y1+y2)/2. Substitute the values and get the answers.

OpenStudy (xoterrixo13):

confused

OpenStudy (tyteen4a03):

@xoterrixo13 Substitute d(1, 2) and e(-3,6) into these 2 equations I just gave you.

OpenStudy (xoterrixo13):

-2,2

OpenStudy (xoterrixo13):

5.656..